# A variational formulation of stochastic thermodynamics. Part I: Finite-dimensional systems

### v1

@@ II.4 Thermodynamic consistency
The average total EP in ST (9) has been shown to always satisfy the second law by definition, due to the properties of the KL divergence. The second law can also be derived by applying Jensen’s inequality to the IFT, which, however, holds for any Markovian stochastic dynamics that do not necessarily represent a physical process [1]. Thus, we can conclude that this second law is not a true thermodynamical law, since it is not influenced by the thermodynamic consistency of the model. This has been previously pointed out in the literature by identifying the total EP as an information measure, termed informatic entropy production rate (IEPR), which, although providing a measure of irreversibility, it cannot be associated with thermodynamic quantities such as heat dissipation or the first law [29, 5].
As a consequence, when applying ST to, for instance, active matter systems, terms without a clear thermodynamic interpretation appear in the EP, identified as unconventional EP [41, 34, 26, 42]. This is because such models, of phenomenological origin, do not explicitly account for the driving mechanisms, therefore lacking thermodynamic consistency [43, 5]. Moreover, the lack of thermodynamic consistency can also be found in models that assume the Einstein relation without taking into consideration multiplicative noise or the associated noise-induced drifts, as it was pointed out in Refs. [14, 5].
Thermodynamic consistency finds different definitions in the literature. In this work, we combine its different meanings and refer to a thermodynamically consistent model as that which verifies:
@@
    Non-decreasing average total EPR $(\dot{S}_{tot}\geq 0)$ [44, 7, 39].
@@
    The generalized Clausius relation Eq. (33) such that the thermodynamic entropy is well-defined [41, 26, 5, 6, 30].
@@
In the following section, we present a variational formulation of ST which systematically incorporates this notion of thermodynamically consistent modeling.
@@ III.1 Thermodynamic phase space
Following the approach described in Ref. [7], we derive the dynamical equations (ODEs) for both configurational and thermal variables by extending the phase space $\Omega$ to include all relevant degrees of freedom. The fundamental thermal variable introduced in this framework is the stochastic thermodynamic entropy $s$. For example, in a thermomechanical system, the extended phase space becomes $\Omega:=\{(\mathbf{x},\mathbf{p},s)\}$.
Hence, $s$ is an independent state variable that accounts for the hidden DoFs of the system and is no longer given as a function of $\mathcal{P}$ as in (12), nor is it given as a function of state [24]. This aligns with the fundamental path-dependence of the medium entropy highlighted in Sec. II.2. Furthermore, this treatment has proven particularly effective in the deterministic (macroscopic) setting for describing nonequilibrium thermodynamics [7], where defining an entropy function is especially challenging [46].
In stochastic descriptions—such as Langevin dynamics—the assumption of time-scale separation renders the heat bath concept intrinsic to the formulation. The thermodynamic entropy $s$ thereby characterizes the macrostate of the bath, whose dynamics are governed by heat exchange with the system, Tds=dqT\,ds=-dq. This interpretation is consistent with the microscopic dynamical entropy framework introduced in Ref. [47], in which the bath contribution to the total entropy can be inferred from the system’s observable DoFs via the Clausius relation.
In contrast to the standard framework of stochastic thermodynamics, the PDF is now defined over the complete thermodynamic phase space $\Omega$, acquiring the form $\mathcal{P}(\mathbf{x},\mathbf{p},s,t)$ following the example of a thermomechanical system. The associated probability measure becomes $\mathcal{P}(\mathbf{x},\mathbf{p},s,t)d\mathbf{x}d\mathbf{p}ds$. Accordingly, the stochastic system entropy and the Shannon entropy are defined as follows:
@@
More generally, we denote the thermal variables by $\varphi_{\alpha}$, which encompass both observable macroscopic quantities—such as the mass $N$ of a chemical species—as well as thermodynamic entropy $s$. These variables span the thermodynamic phase space, which we define as $\Omega:=\{(\mathbf{x},\mathbf{p},\varphi_{\alpha})\}=\{(\bm{\omega},s)\}$, where $\bm{\omega}$ collectively denotes the observable DoFs. Associated with each phase space variable is a conjugate variable, referred to as the (generalized) thermodynamic displacement $\Lambda^{\alpha}$. These displacements play a central role in our formulation, as they enable the derivation of the general form of the dynamical equations through a variational principle.
An additional central quantity in this formulation is the medium entropy, denoted by $\Sigma$. This quantity plays a role analogous to the medium entropy $s_{m}$ in conventional ST, in that it accounts for the entropy production associated with the medium—as defined in Sec. II.3. Introducing $\Sigma$ allows us to distinguish between entropy changes that contribute directly to the system’s thermodynamic entropy $s$ and those arising from interactions with the external environment. The dynamics associated with $\Sigma$ are determined by the nonlinear nonholonomic constraints, formulated within the framework of the Lagrange-d’Alembert principle.
@@ III.2 Lagrange-d’Alembert principle
This internal energy encodes the properties of the heat bath, such as its temperature $T(\mathbf{x},\varphi_{\alpha})\coloneq\partial U/\partial s$ or the exerted force on the system $\mathbf{f}(\mathbf{x},\varphi_{\alpha})\coloneq-\partial U/\partial\mathbf{x}$. These relations correspond the total differential of the internal energy, expressed in terms of the entropy and the observable variables $\bm{\omega}$ as $dU=TdS+\nabla_{\bm{\omega}}U\cdot d\bm{\omega}$ [49]. Notably, this definition of temperature arises explicitly from the internal energy and will appear consistently in the fluctuation-dissipation relation. This differs from earlier instances where $T$ appeared merely as a positive parameter enforcing $T\underline{\underline{\bm{\gamma}}}=\underline{\underline{\mathbf{D}}}$, and was only a posteriori interpreted as the bath temperature.
@@
However, to consistently recover the reversible dynamics, the variational principle must yield conservation laws for the thermal variables in the absence of irreversible processes, i.e., $\dot{\varphi}_{\alpha}=0$. This is accomplished by coupling each thermal variable $\varphi_{\alpha}$ to its conjugate displacement variable $\Lambda^{\alpha}$ within the action functional, through a term of the form $\int dt\,\varphi_{\alpha}\dot{\Lambda}^{\alpha}$ [7]. In the case of entropy, where the time evolution of the medium entropy $\Sigma$ is governed by the nonholonomic constraints, this coupling takes the form $\int dt\,(s-\Sigma)\dot{\Gamma}$, where $\Gamma$ is the conjugate variable to the entropy $s$. The minus sign in this expression ensures that, in the absence of external fluxes, the variational principle gives $\dot{\Sigma}=\dot{s}$,—as expected for an adiabatically closed system—where the thermodynamic EP accounts for the entire medium entropy.
@@
With this preamble, we have shown how irreversible processes can be incorporated within the same geometric structure, with entropy production providing the underlying organizing principle. We then embed this structure into a variational formulation by employing the Lagrange-d’Alembert principle [48, 7], in which the critical curve condition is subject to two constraints: a kinematic constraint on the solution curve and a variational constraint on the variations to be considered when computing the criticality condition. By construction, these constraints encode the information regarding the entropy production. Thus, we define the kinematic and variational nonlinear nonholonomic constraints, respectively, as:
@@
$\mathcal{I}_{p}\coloneq\frac{1}{2}\left[\nabla_{\Omega}\cdot\bm{\mathfrak{F}}-\nabla_{\hat{\Omega}}\cdot\hat{\bm{\mathfrak{F}}}\right$], (39)
and $\bm{\mathfrak{F}}$ collectively denotes all the deterministic forces or fluxes acting on the system (recall $\hat{\Box}$ denotes time-reversal operation). Equation (39) is just the extension of Eq. (32) to the thermodynamic phase space. Since entropy pumping arises from non-dissipative mechanisms, it is not constrained by the FDT and contributes purely through deterministic terms (i.e., $\xi_{\beta}=0$). This is consistent with the standard framework of ST, as such contributions do not explicitly contribute to the total EP, see Eq. (23). A concrete example will be provided in Sec. III.5. See Fig. 1 for an illustration of the different thermodynamic fluxes.
@@
Up to this point, we have not specified any FDRs. The FDRs are obtained by treating the second law as an axiom, imposed on the dynamical equations derived from the variational formulation. Within this framework, the second law is stated as:
$\dot{S}_{tot}(t)\coloneq\dot{\mathcal{S}}(t)+\langle\!\langle\dot{\Sigma}(t)\rangle\!\rangle\geq 0,$ (40)
@@
where $\mathcal{S}(t)$ stands for the Shannon entropy as given by Eq. (34). This definition has the same form as that of ST, i.e., Eq. (22), except that the inequality is only satisfied under the FDRs and not otherwise, contrary to the standard definition in ST. Moreover, this expression is defined in the complete thermodynamic phase space $\Omega$.
In subsequent examples we show how, by construction, this is the only condition we need to impose in this formulation to achieve thermodynamic consistency as defined in Sec. II.4, the rest of conditions arising as a consequence of the second law. Hence, we do not require a specific steady state condition nor the system being either closed or open, such that no thermodynamic equilibrium condition is enforced.
@@
We analyze three different examples: a thermomechanical isolated system (III.3), an interconnected system of multiple reservoirs (III.4), and a thermomechanical open system (III.5).
III.3 Mechanical isolated system

### v2

@@ II.4 Thermodynamic consistency
The average total EP in ST (9) has been shown to always satisfy the second law by definition, due to the properties of the KL divergence. The second law can also be derived by applying Jensen’s inequality to the IFT, which, however, holds for any Markovian stochastic dynamics that do not necessarily represent a physical process [79]. Thus, we can conclude that this second law is not a true thermodynamical law, since it is not influenced by the thermodynamic consistency of the model. This has been previously pointed out in the literature by identifying the total EP as an information measure, termed informatic entropy production rate (IEPR), which, although providing a measure of irreversibility, it cannot be associated with thermodynamic quantities such as heat dissipation or the first law [35, 12, 62].
On a deeper level, this issue stems from the definition of medium entropy itself in ST. Recall the requirement of LDB is expressed as the condition for all paths $\mathbf{z}(t)$ [62]:
$\ln\frac{P[\mathbf{z}|\mathbf{z}_{0}]}{P[\hat{\mathbf{z}}|\mathbf{z}_{t}]}=-\frac{\Delta q}{T}-\Delta\mathcal{I},$ (34)
where the heat bath entropy change is defined as the heat exchanged with the bath divided by its temperature. This notion of entropy has a clear energetic interpretation and coincides with the thermodynamic entropy associated with the bath’s Boltzmann entropy change [93]. This relation holds for any model satisfying LDB [62], for both continuous- and discrete-time dynamics (see e.g. [17]), and therefore inherits the standard properties of thermodynamic consistency. However, in ST the medium entropy is of this form by definition [Eq. (11b)], and hence it is blind to LDB.
As a consequence, when applying ST to, for instance, active matter or biological systems, terms without a clear thermodynamic interpretation appear in the EP, identified as unconventional EP [81, 16, 55, 15]. This is because such models, of phenomenological origin, do not explicitly account for the driving mechanisms. If not all thermodynamically relevant DoFs are taken into account, LDB is violated, the FDRs break down, and entropy production becomes ill-defined due to inconsistent coarse-graining [5, 10, 12]. Moreover, the lack of thermodynamic consistency can also be found in models that assume the Einstein relation without taking into consideration multiplicative noise or the associated noise-induced drifts, as it was pointed out in Refs. [57, 12].
Given the requirement of LDB [Eq. (34)], we refer to a thermodynamically consistent model as that which verifies:
@@
    Non-decreasing average total EPR $(\dot{S}_{\rm tot}\geq 0)$ [29, 44, 32].
@@
    The generalized Clausius relation Eq. (33) such that the thermodynamic entropy has a well-defined physical interpretation [81, 55, 12, 64, 47].
@@
In the following section, we present a variational formulation of (Langevin) ST as a new modeling paradigm which systematically assesses thermodynamic consistency, reconciling the information and physical perspectives of ST.
@@ III.1 Thermodynamic phase space
Following the approach described in Ref. [44], we derive the dynamical equations (ODEs) for both configurational and thermal variables by extending the phase space $\Omega$ to include all thermodynamically relevant DoFs. The fundamental thermal variable introduced in this framework is the stochastic thermodynamic entropy $s$. For example, in a thermomechanical system, the extended phase space becomes $\Omega:=\{(\mathbf{x},\mathbf{p},s)\}$.
Hence, $s$ is an independent variable that accounts for the (internal) hidden DoFs of the system and is no longer given as a function of $\mathcal{P}$ as in (12). Aligning with the fundamental path-dependence of the medium entropy highlighted in Sec. II.2, $s$ is not assumed to be a function of state [89]; path-independence is recovered near equilibrium, as demonstrated in subsequent sections. This treatment has proven particularly effective in the deterministic (macroscopic) setting for describing nonequilibrium thermodynamics [44], where defining an entropy function is especially challenging [34].
In stochastic descriptions—such as Langevin dynamics—the assumption of time-scale separation renders the heat bath concept intrinsic to the formulation. The thermodynamic entropy $s$ thereby characterizes the macrostate of the bath, whose dynamics are governed by heat exchange with the system, Tds=dqT\,ds=-dq. This interpretation is consistent with the microscopic dynamical entropy framework introduced in Ref. [27], where the bath contribution to the total entropy can be inferred from the system’s observable DoFs via the Clausius relation.
In contrast to the standard framework of ST, the PDF is now defined over the complete thermodynamic phase space $\Omega$, acquiring the form $\mathcal{P}(\mathbf{x},\mathbf{p},s,t)$ following the example of a thermomechanical system. The associated probability measure becomes $\mathcal{P}(\mathbf{x},\mathbf{p},s,t)d\mathbf{x}d\mathbf{p}ds$. Accordingly, the stochastic system entropy and the Shannon entropy are defined as follows:
@@
More generally, we denote the thermal variables by $\varphi_{\alpha}$, which encompass both observable macroscopic quantities—such as the mass $N$ of a chemical species—as well as thermodynamic entropy $s$. These variables span the thermodynamic phase space, which we define as $\Omega:=\{(\mathbf{x},\mathbf{p},\varphi_{\alpha})\}=\{(\boldsymbol{\omega},s)\}$, where $\boldsymbol{\omega}$ collectively denotes the observable DoFs. Associated with each phase space variable is a conjugate variable, referred to as the (generalized) thermodynamic displacement $\Lambda^{\alpha}$. These displacements play a central role in our formulation, as they enable the derivation of the general form of the dynamical equations through a variational principle.
An additional central quantity in this formulation is the medium entropy, denoted by $\Sigma$. This quantity plays a role analogous to the medium entropy $s_{m}$ in conventional ST, in that it accounts for the entropy production associated with the medium—as defined in Sec. II.3. Introducing $\Sigma$ allows us to distinguish unambiguously between entropy changes that contribute directly to the system’s thermodynamic entropy $s$ and those arising from interactions with the external environment. The dynamics associated with $\Sigma$ are determined by the nonlinear nonholonomic constraints, formulated within the framework of the Lagrange-d’Alembert principle.
@@ III.2 Lagrange-d’Alembert principle
This internal energy encodes the properties of the heat bath, such as its temperature $T(\mathbf{x},\varphi_{\alpha}):=\partial U/\partial s$ or the exerted force on the system $\mathbf{f}(\mathbf{x},\varphi_{\alpha}):=-\partial U/\partial\mathbf{x}$. These relations correspond the total differential of the internal energy, expressed in terms of the entropy and the observable variables $\bm{\omega}$ as $dU=TdS+\nabla_{\bm{\omega}}U\cdot d\bm{\omega}$ [14]. Notably, this definition of temperature arises explicitly from the internal energy and will appear consistently in the FDR. This differs from earlier instances where $T$ appeared merely as a positive parameter enforcing $T\underline{\underline{\boldsymbol{\gamma}}}=\underline{\underline{\mathbf{D}}}$, and was only a posteriori interpreted as the bath temperature.
@@
However, to consistently recover the reversible dynamics, the variational principle must yield conservation laws for the thermal variables in the absence of irreversible processes, i.e., $\dot{\varphi}_{\alpha}=0$. This is accomplished by coupling each thermal variable $\varphi_{\alpha}$ to its conjugate displacement variable $\Lambda^{\alpha}$ within the action functional, through a term of the form $\int dt\,\varphi_{\alpha}\dot{\Lambda}^{\alpha}$ [44]. In the case of entropy, where the time evolution of the medium entropy $\Sigma$ is governed by the nonholonomic constraints, this coupling takes the form $\int dt\,(s-\Sigma)\dot{\Gamma}$, where $\Gamma$ is the conjugate variable to the entropy $s$. The minus sign in this expression ensures that, in the absence of external fluxes, the variational principle gives $\dot{\Sigma}=\dot{s}$,—as expected for a closed system—where the thermodynamic EP accounts for the entire medium entropy. We recall that $s$ corresponds to the entropy production due to heat dissipation, and is distinct from the system entropy $s_{sys}$ associated with the Shannon entropy.
@@
With this preamble, we have shown how irreversible processes can be incorporated within the same geometric structure, with entropy production providing the underlying organizing principle. We then embed this structure into a variational formulation by employing the Lagrange-d’Alembert principle [44], in which the critical curve condition is subject to two constraints: a kinematic constraint on the solution curve and a variational constraint on the variations to be considered when computing the criticality condition. By construction, these constraints encode the information regarding the entropy production. Thus, we define the kinematic and variational nonlinear nonholonomic constraints, respectively, as:
@@
$\mathcal{I}_{p}:=\frac{1}{2}\left[\nabla_{\Omega}\cdot\bm{\mathfrak{F}}-\nabla_{\hat{\Omega}}\cdot\hat{\bm{\mathfrak{F}}}\right$], (40)
and $\bm{\mathfrak{F}}$ collectively denotes all the deterministic forces or fluxes acting on the system (recall $\hat{\Box}$ denotes time-reversal operation). Equation (40) is just the extension of Eq. (32) to the thermodynamic phase space. Since entropy pumping arises from non-dissipative mechanisms, it is not constrained by the FDT and contributes purely through deterministic terms (i.e., $\xi_{\beta}=0$). This is consistent with the standard framework of ST, as such contributions do not explicitly contribute to the total EP, see Eq. (23). A concrete example will be provided in Sec. III.6. See Fig. 1 for an illustration of the different thermodynamic fluxes.
@@
Let us summarize and provide more clear insight into the variational structure presented here. On one hand we have defined a thermodynamic phase space on which the system Lagrangian takes values; this allows for a variational treatment of thermal variables, as well as a consistent definition of thermodynamic affinities. The critical curve condition [Eq. (III.2)] has been extended to reproduce conservation laws associated with macroscopic variables. Moreover, we have shown the nonholonomic constraints role is manifold: (i) it allows introducing irreversibility in the equations of motion from the variational point of view; (ii) it does so with a clear geometric structure of contraction between fluxes $J_{\gamma}\in\Omega$ and conjugate displacements $\Lambda^{\gamma}\in\Omega^{*}$ $\gamma=(\alpha,\beta)]$, desirable for a rigorous mathematical setting; (iii) such structure restricts the space of all possible trajectories and variations to the set that have a well-defined energetic interpretation of the associated entropy production.
Beyond these geometrical aspects, from the physical point of view, the form of the constraints [Eq. (39)] define a quantity $\Sigma$ which encodes the ratio of stochastic work done by irreversible forces (and the possible entropy fluxes) and bath temperature, which is just the extended Clausius definition of entropy in ST [Eq. (33)], as discussed in Sec. II.3. Such a definition of $\Sigma$ already implies the mentioned assumptions of time-scale separation, large bath size and micro-reversibility, as required by the stochastic description, the Clausius relation and the well-defined energetics associated with entropy production; see Ref. [93] for a detailed discussion on the microcanonical ensemble foundations of ST. Crucially, the constraints define $\Sigma$ and not $s$; recall this allows for an unambiguous distinction between thermodynamic entropy production and external entropy flows. This does not imply that the dynamics of $s$ are left unspecified; owing to the critical curve condition [Eq. (III.2)], the equations of motion for all variables are derived in a self-contained way.
III.3 LDB and the second law
Up to this point, we have not specified the form of irreversible processes. These are determined by imposing the second law, which automatically determines the FDRs linking the fluxes $\{J_{\gamma}\}$ with the noise terms $\{\xi_{\gamma}\}$. As it was argued in Sec. II.4, the second law of ST is blind to thermodynamic consistency (or LDB). However, since this framework provides with a well-defined physical medium entropy $\Sigma$ consistent with micro-reversibility, the condition of LDB implies reconciling the physical definition of $\Sigma$ with the information-theoretic medium entropy $\mathfrak{s}_{m}$. We denote the (information-theoretic) entropy defined in the extended phase space $\Omega$ as $\mathfrak{s}_{m},\mathfrak{s}_{\rm tot}$ in order to distinguish it from $s_{m},s_{\rm tot}$ in the restricted phase space of Sec. II.2, respectively. That is,
$\langle\Delta\mathfrak{s}_{\rm tot}\rangle:=D_{KL}[P[\mathbf{u}]||P[\hat{\mathbf{u}}$]], (41a)
$\Delta\mathfrak{s}_{m}=\ln\frac{P[\mathbf{u}|\mathbf{u}_{0}]}{P[\hat{\mathbf{u}}|\mathbf{u}_{t}]},$ (41b)
for a trajectory $\mathbf{u}:[0,t]\subset\mathbb{R}\rightarrow\Omega$. Recall the system entropy in $\Omega$ was defined in Eq. (35).
@@ III.3 LDB and the second law
This can be achieved by defining the second law as a combination of the information $(\mathcal{S})$ and physical $(\Sigma)$ entropy contributions, imposed as an axiom on the dynamical equations derived from the variational formulation. Within this framework, the second law is then stated as:
$\dot{S}_{\rm tot}(t):=\dot{\mathcal{S}}(t)+\langle\!\langle\dot{\Sigma}(t)\rangle\!\rangle\geq 0,$ (42)
@@
where $\mathcal{S}(t)$ stands for the Shannon entropy as given by Eq. (35). This definition has the same form as that of ST, i.e., Eq. (22), except that the inequality is not inherently satisfied by definition of $\dot{S}_{\rm tot}$, contrary to the standard definition in ST as given in Eq. (23). Moreover, this expression is defined in the complete thermodynamic phase space $\Omega$.
To illustrate this result, recall from Sec. II.2 that $\dot{\mathcal{S}}(t)$ provides with the information-theoretic contributions [Eq. (22)], which in the extended phase space $\Omega$ is denoted as
$\langle\!\langle\dot{\mathfrak{s}}_{\rm tot}\rangle\!\rangle=\dot{\mathcal{S}}+\langle\!\langle\dot{\mathfrak{s}}_{m}\rangle\!\rangle.$ (43)
Although $\dot{\mathfrak{s}}_{m}$ requires path probabilities for its definition and computation at the individual-trajectory level, its ensemble average $\langle\!\langle\dot{\mathfrak{s}}_{m}\rangle\!\rangle$ can be computed from $\mathcal{S}$ alone following (43), as shown in (28).
Then, substituting Eq. (43) back in $\dot{S}_{\rm tot}$:
$\dot{S}_{\rm tot}=\dot{\mathcal{S}}+\langle\!\langle\dot{\Sigma}\rangle\!\rangle=\langle\!\langle\dot{\mathfrak{s}}_{\rm tot}\rangle\!\rangle-\langle\!\langle\dot{\mathfrak{s}}_{m}\rangle\!\rangle+\langle\!\langle\dot{\Sigma}\rangle\!\rangle.$ (44)
In subsequent examples it will be shown how compliance with the inequality (42) requires the solution $\dot{\mathfrak{s}}_{m}=\dot{\Sigma}$, or equivalently,
$\Delta{\mathfrak{s}}_{m}:=\ln\frac{P[\mathbf{u}|\mathbf{u}_{0}]}{P[\hat{\mathbf{u}}|\mathbf{u}_{t}]}=\Delta\Sigma,$ (45)
for a trajectory $\mathbf{u}:[0,t]\subset\mathbb{R}\rightarrow\Omega$. A formal argument for this result is given in Appendix B, and subsequent examples demonstrate this explicitly. Establishing a general proof of uniqueness remains an open problem and an interesting direction for future work. Crucially, Eq. (45) determines the FDRs (of the second kind [62]), which is explicitly shown in Secs. III.4III.6. Moreover, this result implies $\dot{S}_{\rm tot}=\langle\!\langle\dot{\mathfrak{s}}_{\rm tot}\rangle\!\rangle$ and hence the information-theoretic and physical descriptions of entropy are reconciled. As a consequence, $s_{\rm tot}$—the individual-trajectory level version of $S_{\rm tot}$ in Eq. (42)—inherits the relations satisfied by $\mathfrak{s}_{\rm tot}$, which is further discussed in Sec. III.6.
Notably, Eq. (45)—an equality—is the true requirement of LDB extended to the thermodynamic phase space $\Omega$, ensuring its solution provides a consistent physical theory rooted in micro-reversibility. Ultimately, this condition follows from the structure of the nonholonomic constraints (39) (i.e., from the definition of $\Sigma$), which encode the fundamental assumptions underlying LDB; see Ref. [93] for a derivation of LDB. Thus, LDB emerges naturally from the second law axiom [Eq. (42)]. We also show through examples how, by construction, this is the only condition required for thermodynamic consistency as defined in Sec. II.4, the rest of conditions arising as a consequence of LDB.
Note that the specific nature of noise correlations has not yet been specified. Throughout this work, $\{\xi_{\gamma}\}$ are assumed to be zero-mean Gaussian white noises to enable analytical tractability, and our discussion of the results is restricted to this setting. Nevertheless, the Clausius definition of entropy should remain valid for more general stochastic processes [62, 51], implying that the structure of the constraints [Eq. (39)] is preserved, comprising a dissipative drift and its associated fluctuations. It remains, however, an open question under which conditions this form may fail or become analytically intractable when enforcing the second law (or characterizing the average EPR), e.g. for arbitrary temporal correlations.
@@
We analyze three different examples: a thermomechanical isolated system (III.4), an interconnected system of multiple reservoirs (III.5), and a thermomechanical open system (III.6).
III.4 Mechanical isolated system